A boat whose speed is 15 km/hr in still water goes 30 km downstream and comes back in a total of 4 hours 30 minutes. The speed of the stream is ___ km/hr.
1
Expert's answer
2020-07-24T10:08:28-0400

"30\/(15+x)+30\/(15\u2212x)=4.5"


"30(15\u2212x)+30(15+x)=4.5\u22c5(225\u2212x^2)"


"900=1012.5\u22124.5x^2"


"4.5x^2=112.5"


"x=5" "km\/hr" Answer.